AN ELECTRONIC VERSION OF THE TRANSMISSION SPECTRUM OF MOLECULAR IODINE, 14500-20000 cm^{−1}

\underline{A. J. ROSS} and H. SALAMI, Laboratoire de Spectrométrie Ionique et Moléculaire (UMR 5579 CNRS), Bâtiment A. Kastler, Université Lyon I, Domaine Scientifique de la Doua, 69622 Villeurbanne Cedex, France.

The absorption spectrum of the B ← X transition of molecular iodine has been recorded at Doppler resolution, in the region 14500 - 20000 cm^{−1}, on a Fourier transform spectrometer. Calibration has been checked against the Doppler-Free High Resolution Spectral Atlas of the Iodine Molecule^{a} across the region common to both spectra. Our version of the iodine atlas, which was originally produced at Orsay,^{b} avoids the need for systematic wavenumber corrections^{c} to shortwavelength peaks, and should be compatible with any software. It has proved a convenient tool in our laboratory, and we hope others may also find it useful. The transmission spectrum is available, on request, as a single ascii file.

aH. Katô and co-authors, published by the Japan Society for the Promotion of Science, (2000)

bS. Gerstenkorn & P. Luc, Atlas du Spectre d’Absorption de la Molécule d’Iode entre 14800-20000 cm^{−1}, ed. CNRS (1978)

cS. Gerstenkorn & P. Luc, Revue de Physique Appliquée \underline{14} 791 (1979)